# GeoPF: Infusing Geometry into Potential Fields for Reactive Planning in Non-trivial Environments

### v1

@@ IV-A Quantitative Analysis
This section provides a quantitative assessment of the performance of GeoPF and a comparison against baseline state-of-the-art field planners in a simulated environment. [^3] For the simulated scenarios, a collision is considered when the distance between the robot and an obstacle is less than $10^{-8}$.
@@
We evaluate the planners under the following scenarios with randomized conditions, and different spherization strategies (optimized to improve PF and CF results). For each scenario, we generate $1000$ cases randomly sampling obstacles. All scenes were defined in a $x,y,z{\in}[{-}0.2,0.2]$ m workspace with initial $[0,-1,0]$ and goal positions $[0,-1,0]$.
  • Line obstacles are generated with multiple randomly generated line obstacles, as seen in Fig. 8. We classify scenarios as “easy” (5–10 lines) or “hard” (10–50 lines);
  • Scenarios with multiple planes are also classified as “easy” or “hard, according to the number of randomly generated planes, 2–8 for the former and 10-40 planes for hard scenarios. For plane-based scenarios, we also consider “longer” scenarios (changing the goal position);
@@
The results with key performance metrics, [^4] The minimum distance is the smallest value across the trajectory, while the average takes all the robot-obstacle distances through the whole trajectory. for all planners, are shown in Table II and further detailed below. For the maze scenario, we only present an illustrate case in Fig. 9.
@@
For Line Obstacle Scenarios, both the CF and PF require approximating each line segment with small spheres, as seen in Fig. 8. While small spheres ($r^{SP}=0.01$) increase accuracy, they inflate the number of obstacles (and thus computational time). Larger spheres $r^{SP}=0.05$ reduce the count but degrade fidelity. [^5] Since PF is not as competitive as CF [4], we only consider $r^{SP}{=}0.01$. The field parameter for GeoPF was set to $k^{Geo}=0.1$ across all tasks. For PF and CF, we primarily used $k^{attr}=1$, although in some cases, a lower value of $k^{attr}=0.1$ led to improved performance. For each task, we performed a limited parameter search to identify optimal values for PF and CF.
The results, in Table. II, show that GeoPF consistently obtains higher success rates with shorter computational times (up to 8×8$\times$8 × faster). This is partly due to the closed-form distance/force computations for line segments. In contrast, CF and PF often suffer from large overheads due to sphere decompositions, especially in cluttered scenarios.
((a)) Comparison of GeoPF and CF Spherization in Easy (Left) and Hard (Right) Scenarios. The CF parameters in the plot are the one achieves $100\%$ success rate in Table. II First Part. Black lines represent line obstacles, while black circles indicate the spherization of the lines. In cluttered environments, spherization forces the trajectory to navigate around all obstacles. GeoPF plans the trajectory through narrow spaces between obstacles and successfully reaches the goal.
((b)) Spherization of Line and Plane
((c)) Trajectory in Plane Scenario
Figure 8: Comparison of GeoPF and CF Spherization in (a) Ramdom Line and (c) Random Plane Scenarios. (b) illustrates the increased scale of objects (spheres) in spherization from lines to planes, resulting in a sharp increase in computational time. (c) While GeoPF gets a feasible path through obstacles, CF and PF are trapped in local minima, leading to a lower success rate compared to GeoPF in Plane Scenarios.
@@
TABLE II: Quantitative Analysis and comparison between CF, PF, and GeoPF planners. Key metrics: Obs.N. (Number of obstacles), SR (success rate), CT (s) (total computation time), Step (iterations), CT/Step (s), PL (path length), Min and Av. Dist. (distance to obstacles).
Mean / STD M. ($r^{SP},k^{attr}$) Obs.N. SR CT (s) Step CT/Step (ms) PL (m) Min. Dist. (m) Av. Dist. (m)
@@
Mean / STD M. ($r^{SP},k^{SP}$) Obs.N. SR CT (s) Step CT/Step (ms) PL (m) Min. Dist. (m) Av. Dist. (m)
@@
TABLE III: GeoPF performance in complex and dynamic scenarios. Average (black) and standard deviation (gray) shown.
Scenario Obs.N. SR CT (s) Min. Dist. (m)
Complex $12$/$2.45$ 93.7% 3.56/1.37 .032.032.032.032/.03.03.03.03
Dynamic (easy) $\frac{15}{2}(\frac{5}{2})$/$1.1(1.05)$ 94.9% 2.47/0.77 .026.026.026.026/.02.02.02.02
Dynamic (hard) $15(7)$/$1.73(1)$ 83.2% 4.78/1.17 .017.017.017.017/.02.02.02.02
For Plane-based Obstacle Scenarios, we conducted a new parameter search to identify optimal values for PF and CF, as the previously used parameters consistently led to failure. In this evaluation, we analyzed different repulsive coefficients for PF and CF, $k^{SP}=1,10,100$. GeoPF uses the same parameters as in the line task, with $k^{attr}=1$, $k^{Geo}=0.1$.
The results indicate that plane spherization generates a significantly large number of spheres, resulting in an extremely high computational time. Indeed, except for simple line-based scenes, CF and PF planners fail to generate a feasible solution within the real-time framework, as seen in the CT/Step (ms) column in Table II.
In contrast, GeoPF drastically lowers the total computational time—up to 100×100$\times$100 × faster in almost all cases—thanks to direct distance-to-plane calculations. It is also able to keep the real-time performance (computational below 1 ms) for all cases,11 while maintaining a higher success rate in more cluttered arrangements.
The classic Maze Task scenario, where the robot must find a feasible path across multiple planes positioned either vertically or horizontally, is considered a standard trap case. As shown in Fig. 9, while CF and PF fail at the centre point in front of the middle plane, GeoPF successfully navigates through the maze. This result reinforces the findings in Table II which show that CF and PF exhibit significantly lower success in complex scenarios.
@@
We additionally tested GeoPF in challenging environments composed of multiple geometric primitives—including lines, planes, cubes, and cylinders—all placed or moving in random configurations. In these trials, only GeoPF was evaluated, as our goal was to demonstrate its versatility in cluttered, complex, and dynamic scenarios.
Specifically, each complex composite scene randomly selected 5–10 lines, 2–5 planes, and 2–3 cubes or cylinders distributed around the workspace. Some obstacles were static, while others followed randomized trajectories (e.g., drifting planes or cylinders). In Table II, the notation Obs. N. is represented as $a(b)$, where $a$ denotes the total number of random objects, while (b) the dynamic ones, including all geometric primitives.
Throughout these tests, GeoPF maintained high success rates without requiring special tuning or re-initialization. The same $k$ value parameters were used. These results confirm the flexibility and easiness to use, as well as the reliability of GeoPF which preserves its real-time efficiency in more complex scenes with a stable and reliable behaviour.
Figure 10: Non-stationary Scenario. The orange plane moves upward over time. The trajectory passes through the initial plane position as shown in the dark orange area. However, since the plane moves upward along the trajectory evolution, it is no longer an obstacle between robot and goal towards the end. Consequently, the robot follows a (straight) line trajectory in last period of the task.
@@ IV-B Real Robot Experiment
To validate the practical viability of GeoPF, we deployed it on a Franka Emika 7-DoF robotic arm. End-effector velocities were generated by mapping the resultant GeoPF forces into joint-space commands using a standard Jacobian pseudoinverse. Obstacles were tracked via a motion-capture system (Nokov Mars) and fitted to corresponding geometric primitives.
GeoPF was evaluated across three representative tasks. First (Fig. 11(a)), a tabletop scenario, where the robot moved from one corner of the table to another while avoiding randomly placed objects. By leveraging the geometric shapes, GeoPF produced smooth, collision-free trajectories under rigid constraints (fixed z-axis), well-suited for control-based execution. The second task (Fig. 11(c)) involved a 3D setting, navigating around a shelf decomposed into multiple planar elements, along with additional obstacles. GeoPF successfully avoided all obstacles by generating environment-aware trajectories. Lastly, in a dynamic scene (Fig. 1), the robot must pour tea into a moving cup while avoiding randomly moving obstacles. Here, GeoPF updated repulsive forces in real time, consistently maintaining safe distances without parameter adjustment. Indeed, across all experiments, GeoPF generated stable, smooth trajectories within real-time cycles without tuning between scenarios.
@@
Figure 11: Real world experiments. The robot goes from left (low-transparent position) to right (high-opacity) and then returns as seen in trajectories (b),(d). The upper figures show a 2D tabletop task with multiple obstacles, as lines, cubes, and cylinders. The workspace limitation is shown by planes, defining the robot’s boundaries. The bottom figures illustrate a 3D task with four lids representing walls that the robot must avoid during navigation.
@@ V Conclusion
This work introduces a novel reactive planner (GeoPF) that unifies spatial geometric primitives under a temporal potential-field framework for real-time planning. By explicitly leveraging closed-form distance functions for spheres, lines, planes, cubes, and cylinders, GeoPF delivers substantial improvements over existing potential-field methods in computational efficiency, tuning simplicity, and collision avoidance performance. Indeed, quantitative analysis and experimental results demonstrate that GeoPF consistently outperforms existing potential-field-based methods in terms of reliability, ease of deployment, and computational complexity—being the only method suitable for real-time ($1$ kHz) performance in cluttered, unstructured dynamic settings. Future works will involve the integration of prediction modules for obstacle tracking and multi-agent vector fields for effective environment exploration.

### v2

@@ IV-A Quantitative Analysis
We evaluate planners under $1000$ randomized cases per scene within a workspace $x,y,z{\in}[{-}0.2,0.2]$ m with initial position $[0,1,0]$ and goal $[0,-1,0]$. State-of-the-art Baseline planners.
@@
Tuning and spherization values were optimized to improve PF and CF results, under the following designed scenes.
  • Line obstacles: randomly generated lines (Fig. 8) classified as as “easy” (5–10 lines) or “hard” (10–50 lines);
  • Plane obstacles: randomly generated planes, classified as as easy (2-8 planes) or hard (10-40 planes). We also consider “longer” scenarios (changing the goal position);
@@
The results with key performance metrics [^4] The minimum distance is the smallest value across the trajectory, while the average takes all the robot-obstacle distances through the whole trajectory. are shown in Table II.
@@
For Line Obstacle Scenarios, the CF and PF require approximating each segment with small spheres (Fig. 8). While small spheres ($r^{SP}=0.01$) increase accuracy, they greatly increase the number of obstacle primitives and computation time. Larger spheres $r^{SP}=0.05$ reduce complexity but degrade fidelity. [^5] Since PF is not as competitive as CF [3], we only consider $r^{SP}{=}0.01$. GeoPf used a fixed $k^{Geo}=0.1$, across all tasks. PF and CF mainly used $k^{attr}=1$, and $k^{attr}=0.1$ occasionally improving performance. For each task, we performed a limited parameter search to identify optimal values for PF and CF.
Table. II shows GeoPF consistently obtains higher success rates and up to 8×8$\times$8 × faster computation. This is thanks not only to its closed-form distance and force computations, but also to its ability to exploit the geometric structure of lines and planes, differentiating and modulating repulsive forces based on interactions with faces, edges, or vertices. Instead, existing methods rely solely on the closes-contact point—except for CF, which uses sphere-based modulation but introduces significant overhead, especially in cluttered scenes.
@@
Figure 10: Dynamic obstacle avoidance with different primitive-specific modulation, in response to sphere (a), cube (b), and cylinder (c).
TABLE II: Quantitative Analysis (mean/Std) between CF, PF, and GeoPF planners. Key metrics: N.Obs.Pr. (number of obstacle primitives), SR (success rate), Step (iterations), CT/Step (s) (computation time per step), PL (path length), Min and Av. Dist. (distance to obstacles).
M. ($r^{SP},k^{attr}$) N.Obs.Pr. SR Step CT/Step (ms) PL (m) Min. Dist. (m) Av. Dist. (m)
@@
M. ($r^{SP},k^{SP}$) N.Obs.Pr. SR Step CT/Step (ms) PL (m) Min. Dist. (m) Av. Dist. (m)
@@
TABLE III: GeoPF performance in complex and dynamic scenarios. Average (black) and standard deviation (gray) shown. CT/S. represents CT/Step. Min.D. represents Min.Dist.
Scenario N.Obs.Pr. SR CT/S. (ms) Min. D. (m)
Complex $12$/$2.45$ 93.7% $0.62$/$0.37$ $0.032$/$0.03$
Dynamic (easy) $\frac{15}{2}(\frac{5}{2})$/$1.1(1.05)$ 94.9% $0.48$/$0.07$ $0.026$/$0.02$
Dynamic (hard) $15(7)$/$1.73(1)$ 83.2% $0.89$/$0.25$ $0.017$/$0.02$
For Plane-Obstacle Scenarios, [^6] New parameter search was required for PF and CF, as prior values (used successfully in line-based scenes) consistently failed. We analyzed repulsive coefficients $k^{SP}=1,10,100$. GeoPF kept the same parameters. spherization led to massive sphere counts and excessive computational times. Except in simple cases, PF and CF could not produce feasible solutions within real-time constraints, as shown in Table II. In contrast, GeoPF reduced computational time by up to 100×100$\times$100 ×, achieving real-time (below 1 ms) in all cases while maintaining higher success rates even in dense and cluttered scenarios.
In the challenging Maze Task, considered a standard trap case for reactive planners with multiple vertical and horizontal planes (Fig. 9), CF and PF fail at the centre point in front of the middle plane, GeoPF successfully navigates through the maze, aligning with results from Table II.
@@
We further evaluated in challenging environments composed of multiple primitives—including lines, planes, cubes, and cylinders, in cluttered and dynamic configurations. Each composite scene randomly included 5–10 lines, 2–5 planes, and 2–3 cubes or cylinders distributed around the workspace.
GeoPF consistently achieved high success rates without requiring special tuning or re-initialization, see Table III. Fig. 10 further shows the dynamic response modulated per primitive in a back-and-forth scene (with dynamic obstacles appearing in the middle of an originally straight line goal). The same $k$ value parameters were used throughout this paper. These results highlight GeoPF’s flexibility, real-time performance, and stable behavior in complex, dynamic scenes.
@@ IV-B Real Robot Experiment
To validate GeoPF’s practical viability, we deployed it on a 7-DoF Franka Robot. End-effector velocities were generated by mapping the GeoPF forces into joint-space via Jacobian pseudoinverse. Obstacles were tracked via a motion-capture system (Nokov Mars) and fitted to corresponding geometric primitives. GeoPF was tested across four representative tasks.
First (Fig. 11(a)), in a tabletop scenario, GeoPF leveraged geometric primitives to produce smooth, collision-free trajectories under rigid (z-axis) constraints, avoiding randomly placed lines, cubes, and cylinders within plane-bounded workspaces. To further evaluate robustness to primitive shaping, we revised the same experiment, but using only bounding boxes (from off-the-shelf algorithm, e.g., YOLOv7). As seen in Fig. 12(b), even such coarse approximations suffice for GeoPF to generate safe, efficient and real-time motions in clutter, while state-of-the-art methods Fig. 12(c)-(d) failed despite access to full geometry.
Second, in a 3D task, the robot navigated around a shelf decomposed into multiple planar elements, along with additional obstacles. GeoPF successfully avoided all obstacles by generating environment-aware trajectories.
In the dynamic scene shown in Fig. 1, the robot performed tasks like pouring tea into a moving cup or navigating around moving obstacles, with GeoPF consistently maintaining safe distances in real time without parameter adjustment. Across all cases, GeoPF produced real-time smooth, collision-free trajectories suitable for control-based execution.
Lastly, we tested GeoPF for simultaneous EE and whole-body collision avoidance (Fig. 13). The links were shaped as cylinders, actively modulating reactive forces as detailed in Sec. III-E. The resulting forces were then mapped to torques based on the moment arm to the previous joint. This enabled coordinated, geometry-aware avoidance across the entire kinematic chain without additional tuning.
@@
Figure 11: Real world experiments. The robot goes from left to right (high-opacity) and then returns, see trajectories (b),(d). The upper figures show a 2D tabletop task with multiple obstacles, as lines, cubes, and cylinders. The workspace is shown by planes, w.r.t. the robot’s boundaries. The bottom figures illustrate a 3D task with four lids representing walls that the robot must avoid during navigation.
((a))
((b))
((c))
((d))
Figure 12: Robustness to primitive fitting—based on 2D Task (Fig. 11(a)). (a) GeoPF with structured primitives-fitting produces high-quality, shape-aware trajectories, while CFs (c) and APFs (d) fail to generate a collision-free trajectory, despite being given full geometric-information. (b) GeoPF with only access to bounding boxes still finds a feasible path—even with less privileged information. The red lines indicate that robot goes into the obsctale.
((a)) Reactive (no avoidance)
((b)) GeoPF (w/o whole-body)
((c)) GeoPF w/ whole-body)
Figure 13: Whole-body collision avoidance with GeoPF. (a) Planner with no avoidance, naturally leading to multiple collisions, (b) End-effector-only avoidance fails to avoid elbow collisions. (c) GeoPF’s whole-body geometric-aware collision avoidance successfully modulates arm and end-effector trajectories to safely avoid all obstacles.
@@ V Conclusion
This work introduces GeoPF, a novel reactive planner that unifies spatial structured geometric primitives under a temporal potential-field framework for real-time planning. By explicitly leveraging closed-form distance functions and structural and spatial properties of primitives, GeoPF delivers substantial improvements in computational efficiency, tuning simplicity, and collision avoidance performance. Experiments demonstrate its robustness across static and dynamic scenarios, including whole-body avoidance, with minimal parameter adjustment. Quantitative analysis and experiments demonstrate its reliability, robustness, and easy-deployment over existing field-based planners, being the only suitable for real-time ($1$ kHz) performance in cluttered, unstructured dynamic settings. GeoPF has also been tested for whole-body avoidance and limited primitive-fitting access, all with no tuning.
Future work will focus on integrating prediction modules for dynamic obstacle tracking and multi-agent vector fields to address a core limitation of GeoPF—and all reactive planners—its lack of completeness. We also aim to fuse GeoPF’s primitive structure with neural SDFs (e.g., [27]), extending ANN inference to produce rich geometric descriptors.